OpenStudy (anonymous):

The volume of a pyramid that fits exactly inside a cube is 18 cubic feet. What is the volume of the cube? 6 cubic feet 18 cubic feet 54 cubic feet 72 cubic feet

OpenStudy (anonymous):

@whpalmer4 can you help with this one

OpenStudy (whpalmer4):

So this is a pyramid with a square base, and its height is exactly the same as the length of a side of the base, is that how you see it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

where do we go from there

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

do you remember the formula for the volume of such a pyramid?

OpenStudy (anonymous):

no i dont @whpalmer4

OpenStudy (whpalmer4):

okay, I didn't either, so I looked it up :-) \[V = \frac{1}{3}Ah\]where \(A\) is the area of the base. But as we have a square base, with side length \(s\), and our height is equal to the side length of the base, what is our formula going to be?

OpenStudy (anonymous):

V=1/3 (64)(8) ? @whpalmer4

OpenStudy (whpalmer4):

oh, we don't know any numbers... \[A = s^2\] if the side length is \(s\), right? and \[h = s\] if the height is equal to the side length (it should be, since we are fitting this into a cube where it just fits) so \[V = \frac{1}{3}Ah = \frac{1}{3}s^2*s = \frac{1}{3}s^3\]right?

OpenStudy (anonymous):

im confused, do you mean 18? @whpalmer4

OpenStudy (whpalmer4):

Yeah, that's the next step. I'm just finding the formula for the volume of a rectangular prism that is fit inside a cube, so we could do this problem with any size cube or pyramid we like. So, do you agree with how I derived that formula?

OpenStudy (anonymous):

yes @whpalmer4

OpenStudy (whpalmer4):

Good. that's the important part here. Now, we know that for this particular pyramid, V = 18 ft^3, so \[V = 18 \text{ ft}^3 = \frac{1}{3}{s^3}\]If we rearrange that a bit, we have \[54 \text{ ft}^3 = s^3\] Now, what is the volume of the cube enclosing our pyramid? Remember, we've already decided it has side length \(s\)...

OpenStudy (anonymous):

54? @whpalmer4

OpenStudy (whpalmer4):

No, your answer should be "\(54\text{ ft}^3\)", and said with some certainty :-)

OpenStudy (anonymous):

thanks so much @whpalmer4

OpenStudy (whpalmer4):

Notice that we don't ever have to actually figure out the side length or height, which is good, because it's kind of a clumsy number, \(s = 3\sqrt[3]{2} \approx 3.77976\) We can test our answer, now that I did so: \[V_{cube} = (3\sqrt[3]{2})^3 = 3^3*2 = 27*2 = 54\checkmark\]\[V_{pyramid} = \frac{1}{3}(3\sqrt[3]{2})^2*3\sqrt[3]{2} = \frac{1}{3}*9*3*2 = 18\checkmark\]

OpenStudy (whpalmer4):

Any questions?